Zwee:

Given the functions f(n) = 25 and g(n) = 3(n − 1), combine them to create an arithmetic sequence, an, and solve for the 12th term. (2 points)

jhonyy9:

like a first step you need to know what is an aritmetic sequence - do you know ? @Laylalyssa

jhonyy9:

so - in an arithmetic sequence the difference between one term and the next is a constant @AZ opinon pls. ty

jhonyy9:

how you like - wan combine them ? adding ?

jhonyy9:

so in this way what will get ? f(n) +g(n) = ?

jhonyy9:

= 25 + 3(n-1) = distribute parentheses what will get ?

jhonyy9:

= 25 + 3n -3 = 22 +3n so we get f(n) +g(n) = 22 +3n how you get the first term ? @Laylalyssa

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 = 25 + 3n -3 = 22 +3n so we get f(n) +g(n) = 22 +3n how you get the first term ? @Laylalyssa \(\color{#0cbb34}{\text{End of Quote}}\) I never learned this

jhonyy9:

a_n = 22 +3n a_1 = ?

jhonyy9:

to get a_1 give to n the value of 1

jhonyy9:

a_1 = ?

jhonyy9:

@Zwee

jhonyy9:

a_n = 22 +3n a_1 = 22 +3*1 = ?

jhonyy9:

@Laylalyssa

jhonyy9:

a_1 = 22 +3*1 = 22 +3 = 25 a_2 = 22 +3*2 = ?

jhonyy9:

@supie do you can continue it ?